Fermi National Laboratory


Accelerator Update for June 24, 2005 - June 27, 2005

Friday June 24

The day shift began with Operations monitoring the Recycler stash, with the Antiproton Source (Pbar) stacking, and with the following experiments taking beam: MiniBooNE, and NuMI, and Meson experiments MCenter and MTest.

Operations put the Tevatron (TeV) into a dry squeeze at 8:25 AM.

Operations put the TeV into shot setup at 10:14 AM.

Operations established combination store 4233 at 12:48 PM with an initial luminosity of 101.9E30. Pbar resumed stacking.

At 2:21 PM, a power supply expert began investigating a bad fan on a Main Injector (MI) sextupole (I:SEXVPS). This halted stacking and beam to NuMI. The expert discovered that a fuse had blown, replaced the fuse, and had the supply back up by 2:46 PM.

At 3:13 PM, experts opened the transformer for E:V108 to troubleshoot a problem. NuMI was off during this investigation. The experiment was back on line fifteen minutes later. The experts found a small leak, but said that no immediate action was necessary.

At 6:21 PM, Operations reported that the TeV sector B1 pond water pump had failed. The expert said the pump would be replaced on Saturday.

Operations put the Recycler into shot setup at 6:54 PM to stash antiprotons.

The Recycler stash ended and Pbar resumed stacking at 8:12 PM. MCenter and MTest resumed taking beam at 8:26 PM.

Saturday June 25

The midnight shift began with Operations monitoring store 4233 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, MCenter, and MTest taking beam.

Recycler began stashing antiprotons at 12:24 AM.

MTest experimenters asked to have their beam stopped for the night at 12:40 AM.

The Recycler stash ended and MCenter resumed taking beam at 1:55 AM.

Operations reestablished beam to MTest at 7:58 AM.

At 8:54 AM, a MI RF station (MIRF5) tripped off for the first of 20 times in about three hours. It reset okay each time. An expert was called for advice.

Operations terminated store 4233 at 5:03 PM.

Operation put the TeV into shot setup at 5:16 PM.

Operations established store 4234 at 7:17 PM with an initial luminosity of 115.92E30.

At 7:28 PM, an Accumulator SEPTA (A:ISEP1V) tripped off. Operators managed to get the device back on and resume stacking at 7:46 PM. MCenter and MTest beam resumed a bit later.

Recycler began stashing at 11:40 PM.

Sunday June 26

The midnight shift began with Operations monitoring store 4234, with the Recycler stashing, and with MiniBooNE and NuMI taking beam.

The Recycler stash ended and Pbar resumed stacking at 12:35 AM. MCenter resumed taking beam.

MTest resumed taking beam around 1 AM.

Recycler began stashing around 5:13 AM. The transfers were completed by 5:59 AM. Pbar resumed stacking and MCenter and MTest resumed taking beam.

At 6:54 AM, Operations reported that the Booster RF system had been sparking for about fifteen minutes. BRF15 had tripped off and took awhile to reset properly. Operators investigated and then called in experts.

At 7:23 AM, Operations reported that BRF8 & 15 tripped due to sparking. BRF13 tripped off at 7:39 AM and wouldn’t reset. Operators reported that all this sparking was causing serious beam problems.

At 8:20 AM, Operations reported that MiniBooNE was unable to run due to the Booster problems. Operators turned Booster down to one turn (a very low intensity) and managed to send some beam to MiniBooNE. Experts continued their investigations. By 8:38 AM, operators managed to get the Booster up to eight turns. Operations then shut off the Booster beam for about ten minutes so the experts could take MIRF13 out of the RF system.

At 11:49 AM, the Booster RF experts reported that they needed a tunnel access to make repairs. Operators racked out the power to Booster at 12:12 PM; this shut down all downstream beam.

Booster came out of access at 5:17 PM. The experts repaired and brought all the Booster RF stations back on line.

Pbar resumed stacking at 6:08 PM and MCenter and MTest resumed taking beam. Booster experts began tuning their machine

A TeV expert investigated the store’s luminosity drop. He reported around 8 PM that the problem was due to movement in the TeV sector B0 low beta quadrupoles and there wasn’t much that could be done about it.

Operations terminated store 4234 at 11:03 PM and then put the TeV into shot setup at 11:19 PM.

Monday June 27

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash and stack, and with MiniBooNE and NuMI taking beam.

Operations established store 4235 at 2:49 AM with an initial luminosity of 98.98E30.

Operations resumed sending beam to MCenter at 3:12 AM. MTest gave up for the night.

Pbar didn’t resume stacking until 4:26 AM due to SEPTA problems with the Debuncher and Accumulator.

The Plan for Monday

The plans for today are to stack, stash, and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/27/2005   email Fermilab